A gas mixture containing oxygen, nitrogen, and helium exerts a total pressure of 925 torr .

If the partial pressures are oxygen 445 torr and helium 50 torr, what is the partial pressure (torr) of the nitrogen in the mixture?
1
Expert's answer
2020-10-19T14:20:41-0400

"P=p(O_2)+p(N_2)+p(He)"

"p(N_2)=P-p(O_2)-p(He)"

"p(N_2)=(925-445-50)torr=330torr"

"p(N_2)=330torr"